Jenny wants to utilize a portion of her garden for growing flowers. Being a designer herself, she visualizes a triangular area with different lengths for each side. For this Jenny decides to keep a side of the triangular area 5 feet shorter and the other 4 feet longer than the third side. She also wants to restrict the perimeter of the triangular area to 29 feet, so that it does not cover a huge area of the garden. Determine the length of each side of the triangular area that Jenny has visualized.

Answer by mangopeeler07(462)

side 1=x-5
side 2=x+4
side 3=x

Side 1+ side 2+ side 3=29
(x-5)+(x+4)+(x)=29

Combine like terms
3x-1=29

Add one to each side
3x=30

Divide by 3
x=10

Therefore,
side 1=5
side 2=14
side 3=10